Company D is comparing two different capital structures: an all-equity plan (Plan I) and a levered plan (Plan II). Under Plan I, the company would have 160,000 shares of stock outstanding. Under Plan II, there would be 80,000 shares of stock outstanding and $3million in debt outstanding. The interest rate on the debt is 5%, and there are no taxes. (a)If EBIT is $300,000, which plan will result in the higher EPS? (b)If EBIT is $500,000, which plan will result in the higher EPS? (c)What is the break-even EBIT?
Answer #1
a)
Plan 1
EPS = EBIT*(1-tax rate)/shares = 300000*(1-0)/160000=1.88
Plan 2
EPS = (EBIT-debt*interest rate)*(1-tax rate)/shares = (300000-3000000*0.05)*(1-0)/80000=1.88

both have same EPS

b)
Plan 1
EPS = EBIT*(1-tax rate)/shares = 500000*(1-0)/160000=3.13
Plan 2
EPS = (EBIT-debt*interest rate)*(1-tax rate)/shares = (500000-3000000*0.05)*(1-0)/80000=4.38

Plan 2 has higher EPS

c)Break even EBIT is the EBIT where EPS plan I = EPS plan II
EBIT*(1-tax rate)/shares = (EBIT-interest rate*debt)*(1-tax rate)/shares
EBIT*(1-0)/160000=(EBIT-0.05*3000000)*(1-0)/80000
EBIT =300000